OpenStudy (anonymous):

Complete these ordered pairs for this equation. (0, ), (-2, ), (4, ) y=2x @heather040200

OpenStudy (anonymous):

I believe you plug you do this y=2(1) to get y for the first one y=2(-2) y=2(4)

OpenStudy (anonymous):

(0,1)(-1,-2),(2,4)

OpenStudy (anonymous):

no you have to multiply

OpenStudy (anonymous):

i dont get it i alreday try and i cant get the right answer

OpenStudy (anonymous):

you just have to plug in the numbers you are given for x y=2(0)=0 (0,0) y=2(-2)=-4 (-2,-4) y=2(4)=8 (4,8)

OpenStudy (anonymous):

do you understand??

OpenStudy (anonymous):

kind of i am ganna post another on so you ganna be tag in it

OpenStudy (anonymous):

kkk